OpenStudy (neycuurose1):

Transform this equation to slope-intercept form: 6x + 4y = 8. A. y= 3/2 x +4 B. y= -3/2 x + 2 C. y= 3/4x +2 D. y= -3/2 x + 4 I'm battling between A and B... Yet I'm still unsure.

rebeccaxhawaii (rebeccaxhawaii):

get y by itself

OpenStudy (neycuurose1):

Well, y = -2 + (3/2)x @rebeccaxhawaii

OpenStudy (neycuurose1):

Idk

OpenStudy (neycuurose1):

oh my bad I think I solved for y

OpenStudy (anonymous):

You have to solve for y, if i'm not wrong.

OpenStudy (neycuurose1):

I think you might be right. You have to solve for y first, but I'm so confoosed.

OpenStudy (anonymous):

Ik here are the steps: 6x+4y=8 4y=8-6x y=-(3/2)x+2

OpenStudy (neycuurose1):

thanks, I think I was doing it wrong until I got so many answers o.o

OpenStudy (neycuurose1):

Here's a medal c:

OpenStudy (anonymous):

Thanks! Medal back ;D

OpenStudy (anonymous):

Bye

OpenStudy (anonymous):

Tag me if you need help.